Mrs. Shieh was 24 years old when Ingred was born. In three years the sum of their ages will be 68 years. How old is each now?: Mrs. Shieh was 24 years old when Ingred was born. In three years the sum of their ages will be 68 years. How old is each now?
Answer by ankor@dixie-net.com(4507)

Mrs. Shieh was 24 years old when Ingred was born. In three years the sum of their ages will be 68 years. How old is each now?
:
We can say that S is 24 yrs older than I, therefore
S = (I+24)
:
An equation for the statement:
"In three years the sum of their ages will be 68 years."
(S+3) + (I+3) = 68
S + I + 6 = 68
S + I = 68-6
S + I = 62
:
How old is each now?
:
Substitute (I+24) for S in the above equation
(I+24) + I = 62
2I = 62 - 24
I = 38/2
I = 19 yrs
then
19+24 = 43 yrs is S
:
:
Check solution in the statement:
"In three years the sum of their ages will be 68 years."
(43+3) + (19+3) = 68